# A uniform-in-time nonlocal approximation of the standard Fokker-Planck equation

### v3

@@ 3.3 An $L^{\infty}$ version of Berry-Esseen’s central limit theorem
(NOT THE UPDATED VERSION: $l/L$) First, we notice that $\mathfrak{f}_{n}$ satisfies the hypotheses of Lemma 3.9 for every $n$, and, in particular
$\int_{\mathbb{R}^{d}}\mathfrak{f}_{n}(x)x_{i}x_{k}\,\mathrm{d}x=\delta_{ij}.$
@@
$\begin{split}{\hat{\mathfrak{f}}_{n}}(\xi)&=\int_{\mathbb{R}^{d}}e^{-ix\xi}(\bar{\sigma}^{2}n)^{d/2}f^{\sigma_{1}\dots,\sigma_{n}}(\sqrt{n}\bar{\sigma}x)\,\mathrm{d}x=\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi}{\sqrt{n}\bar{\sigma}}}f_{\sigma_{1}}*\cdots*f_{\sigma_{n}}(y)\,\mathrm{d}x\\ &=\mathcal{F}(f_{\sigma_{1}}*\cdots*f_{\sigma_{n}})\Big(\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big)=\prod_{j=1}^{n}\hat{f}_{\sigma_{j}}\Big(\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big)=\prod_{j=1}^{n}\int_{\mathbb{R}^{d}}e^{-ix\frac{\xi}{\sqrt{n}\bar{\sigma}}}\sigma_{j}^{-d}f\Big(\frac{x}{\sigma_{j}}\Big)\,\mathrm{d}x\\ &=\prod_{j=1}^{n}\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}}f(y)\,\mathrm{d}y=\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big),\end{split}$
@@
$\left\|\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right\|_{L^{1}}\leq\prod_{j=1}^{n}\left\|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right\|_{L^{n}}=\prod_{j=1}^{n}\Big(\frac{\bar{\sigma}\sqrt{n}}{\sigma_{j}}\Big)^{d/n}\left\|\hat{f}\right\|_{L^{n}}\leq n^{d/2}\Big(\frac{L}{l}\Big)^{d}\left\|\hat{f}\right\|_{L^{n}}^{n},$
@@
$\left\|\mathfrak{f}_{n}-G\right\|_{L^{\infty}}\leq\int_{\xi\in B^{c}_{\delta,n}}|\hat{\mathfrak{f}}_{n}|\,\mathrm{d}\xi+\int_{\xi\in B^{c}_{\delta,n}}|\hat{G}|\,\mathrm{d}\xi+\int_{\xi\in B_{\delta,n}}|\hat{\mathfrak{f}}_{n}-\hat{G}|\,\mathrm{d}\xi:=T_{1}+T_{2}+T_{3}$
The first two terms can be bounded in a similar way. For $T_{1}$, and without loss of generality, assume $p^{\prime}$ is an integer (otherwise, we can work with $\left\lceil p^{\prime}\right\rceil$, the smallest integer larger or equal than $p^{\prime}$, since $\hat{f}$ is also in $L^{\left\lceil p^{\prime}\right\rceil}$). Then
$\int_{B_{\delta}^{c}}\prod_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\,\mathrm{d}\xi=n^{d/2}\int_{|\eta|\geq\frac{L}{l}\delta}\prod_{j=1}^{n}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta\\ \leq n^{d/2}\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}+1}}{\bar{\sigma}}\eta\Big)\right|\cdots\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{n}}{\bar{\sigma}}\eta\Big)\right|\int_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\right|\cdots\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta.$ (18)
@@
$\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|\leq\sup_{|\eta|\geq\frac{\bar{\sigma}}{\sigma_{i}}\delta}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|=\sup_{|\eta^{\prime}|\geq\delta}|\hat{f}(\eta^{\prime})|\leq\kappa(\delta)$ (19)
@@
$\begin{split}\int_{|\eta|\geq\frac{L}{l}\delta}&\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\right|\cdots\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta\\ &\leq\bigg(\int_{\mathbb{R}^{d}}\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg)^{1/p^{\prime}}\cdots\bigg(\int_{\mathbb{R}^{d}}\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg)^{1/p^{\prime}}\\ &=\Big(\frac{\sigma_{1}}{\bar{\sigma}}\Big)^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\cdots\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\Big)^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\leq\Big(\frac{L}{l}\Big)^{d}\|\hat{f}\|_{p^{\prime}}^{p^{\prime}}\leq\Big(\frac{L}{l}\Big)^{d}C_{p}\|f\|_{p}^{p^{\prime}},\end{split}$ (20)
@@
We proceed analogously for $T_{2}$, obtaining a very similar bound also for $T_{2}$. Since $\kappa(\delta)<1$, $\kappa(\delta)^{n-p^{\prime}}$ decays very fast in $n$, thus for $n$ large enough (depending on $p$), there exists $C_{1}>0$ such that
@@
$\left|\hat{G}\Big({\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}}\Big)\right|=(2\pi)^{-d/2}e^{-\frac{|\xi|^{2}}{2n}\frac{\sigma_{k}^{2}}{\bar{\sigma}^{2}}}\leq e^{-\frac{|\xi|^{2}}{2n}(\frac{l}{L})^{2}}.$
@@
$\begin{split}\frac{\left|\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}(\xi)\right|}{|\xi|^{2+s}}&=\frac{1}{\left|\xi\right|^{2+s}}\Big[\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=1}^{j-1}\hat{f}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=j+1}^{n}\hat{G}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\Big]\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|e^{-\frac{|\xi|^{2}}{4n}(j-1)(\frac{l}{L})^{2}}e^{-\frac{|\xi|^{2}}{2n}(n-j)(\frac{l}{L})^{2}}\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{j=1}^{n}\left|\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}\frac{\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|}{\left|\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}}e^{-\frac{|\xi|^{2}}{4}\frac{n-1}{n}(\frac{l}{L})^{2}}\\ &\leq\frac{1}{n^{\frac{2+s}{2}}}\Big(\frac{L}{l}\Big)^{2+s}n\sup_{\eta^{\prime}}\frac{|\hat{f}(\eta^{\prime})-\hat{G}(\eta^{\prime})|}{\left|\eta^{\prime}\right|^{2+s}}e^{-\frac{\left|\xi\right|^{2}}{8}(\frac{l}{L})^{2}}.\end{split}$
@@
$\sup_{\eta^{\prime}}\frac{|\hat{f}(\eta^{\prime})-\hat{G}(\eta^{\prime})|}{\left|\eta^{\prime}\right|^{2+s}}\leq\sup_{\eta^{\prime}}\left(\frac{|\hat{f}(\eta^{\prime})-\hat{f}(0)|}{\left|\eta^{\prime}\right|^{2+s}}+\frac{|\hat{G}(\eta^{\prime})-\hat{G}(0)|}{\left|\eta^{\prime}\right|^{2+s}}\right)\leq C_{1}(\rho_{2+s}+M_{2+s}(G)).$
@@ 3.4 Positivity of the solution
the la (not the updated version) Let $X_{m}$ be a Poisson random variable of parameter $m$. For a fixed $m$, using a known result on bounds for the tail probabilities of a Poisson random variable, we have that
$s_{m}=1-\P(X_{m}<m)-\P(X_{m}\leq 2m+1)\geq\frac{1}{2}-e^{m+1}\Big(\frac{m}{2m+1}\Big)^{2m+1}:=b_{m}\geq b_{1}=:C_{L},$
since $b_{m}$ is increasing for $m\geq 1$. To prove that $s_{m}\to 1/2$, let us now take a random variable $X_{m}$, Poisson-distributed with parameter $m$. Let $Y_{i},i=1,\dots,m$ be Poisson-distributed random variables with parameter $1$ so that $X_{m}=\sum_{i=1}^{m}Y_{i}$. We have
$\begin{split}&e^{-m}\sum_{n=m}^{m}\frac{m^{n}}{n!}=e^{-m}\sum_{n=0}^{2m}\frac{m^{n}}{n!}-e^{-m}\sum_{n=0}^{m-1}\frac{m^{n}}{n!}\\ &=\mathbb{P}\Big(\sum_{i=1}^{m}Y_{i}-m\leq{m}\Big)-\mathbb{P}\Big(\frac{1}{\sqrt{m}}\big({\sum_{i=1}^{m}Y_{i}-m}\big)\leq 0\Big)\ \rightarrow 1-\frac{1}{2}=\frac{1}{2},\end{split}$
@@
(put it twice: once after Lyapunov and once here) Theorem 3.2 gives the confining condition on the generator $L_{\varepsilon}$ in $L^{1}_{\left\langle k\right\rangle}$ and consequently in $L^{1}_{k}$. Let $u_{0}$ be in $\mathcal{D}(L_{\varepsilon})$. By standard semigroup theory $u(t)\in\mathcal{C}^{1}([0,\infty);L^{1}_{k})$, and consequently we compute with $\varphi=\left\langle x\right\rangle^{k}$,
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{d}}u(t,\cdot)\varphi(x)\,\mathrm{d}x&=\int_{\mathbb{R}^{d}}\frac{\mathrm{d}}{\mathrm{d}t}u(t,\cdot)\varphi(x)\,\mathrm{d}x\\ =&\int_{\mathbb{R}^{d}}(L_{\varepsilon}u)\varphi(x)\stackrel{{\scriptstyle\mathclap{\mbox{*}}}}{{=}}\int_{\mathbb{R}^{d}}u(L_{\varepsilon}^{*}\varphi)\leq\int_{\mathbb{R}^{d}}u(-\lambda\varphi+C)\end{split}$
[…]
$f^{\prime}(t)\leq-\lambda f(t)+C\implies f(t)\leq e^{-\lambda t}f(0)+\frac{C}{\lambda}(1-e^{-\lambda t})$
Since $\mathcal{D}(L_{\varepsilon})$ is dense in $L^{1}_{k}$, the claim is true also for any $u_{0}\in L^{1}_{k}$ by standard density arguments, hence proving the Lyapunov condition on semigroups. Similarly, part $(ii)$ of Theorem 1.1 is an immediate consequence of Theorem 3.10 and Theorem 3.4, proving a semigroup Lyapunov condition in a similar manner with $\varphi=e^{a\left\langle x\right\rangle}.$

### v4

@@ 3.3 An $L^{\infty}$ version of Berry-Esseen’s central limit theorem
First, we notice that $\mathfrak{f}_{n}$ satisfies the hypotheses of Lemma 3.9 for every $n$, and, in particular
$\int_{\mathbb{R}^{d}}\mathfrak{f}_{n}(x)x_{i}x_{j}\,\mathrm{d}x=\delta_{ij}.$
@@
$\begin{split}{\hat{\mathfrak{f}}_{n}}(\xi)&=\int_{\mathbb{R}^{d}}e^{-ix\xi}(\bar{\sigma}^{2}n)^{d/2}f^{\sigma_{1}\dots,\sigma_{n}}(\sqrt{n}\bar{\sigma}x)\,\mathrm{d}x=\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi}{\sqrt{n}\bar{\sigma}}}f_{\sigma_{1}}*\cdots*f_{\sigma_{n}}(y)\,\mathrm{d}y\\ &=\mathcal{F}(f_{\sigma_{1}}*\cdots*f_{\sigma_{n}})\Big(\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big)=\prod_{j=1}^{n}\hat{f}_{\sigma_{j}}\Big(\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big)=\prod_{j=1}^{n}\int_{\mathbb{R}^{d}}e^{-ix\frac{\xi}{\sqrt{n}\bar{\sigma}}}\sigma_{j}^{-d}f\Big(\frac{x}{\sigma_{j}}\Big)\,\mathrm{d}x\\ &=\prod_{j=1}^{n}\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}}f(y)\,\mathrm{d}y=\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big),\end{split}$
@@
$\left\|\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right\|_{L^{1}}$ $\leq\prod_{j=1}^{n}\left\|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right\|_{L^{n}}=\prod_{j=1}^{n}\Big(\frac{\bar{\sigma}\sqrt{n}}{\sigma_{j}}\Big)^{d/n}\left\|\hat{f}\right\|_{L^{n}}$
$\leq n^{d/2}\Big(\frac{L}{l}\Big)^{d}\left\|\hat{f}\right\|_{L^{n}}^{n},$
@@
$\left\|\mathfrak{f}_{n}-G\right\|_{L^{\infty}}\leq\int_{\xi\in B^{c}_{\delta}}|\hat{\mathfrak{f}}_{n}|\,\mathrm{d}\xi+\int_{\xi\in B^{c}_{\delta}}|\hat{G}|\,\mathrm{d}\xi+\int_{\xi\in B_{\delta}}|\hat{\mathfrak{f}}_{n}-\hat{G}|\,\mathrm{d}\xi:=T_{1}+T_{2}+T_{3}$
The first two terms can be bounded in a similar way. Without loss of generality, assume $p^{\prime}$ is an integer (otherwise, we can work with $\left\lceil p^{\prime}\right\rceil$, the smallest integer larger or equal than $p^{\prime}$, since $\hat{f}$ is also in $L^{\left\lceil p^{\prime}\right\rceil}$). Then
$\int_{B_{\delta}^{c}}\prod_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\,\mathrm{d}\xi=n^{d/2}\int_{|\eta|\geq\frac{l}{L}\delta}\prod_{j=1}^{n}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta\\ \leq n^{d/2}\sup_{|\eta|\geq\frac{l}{L}\delta}\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}+1}}{\bar{\sigma}}\eta\Big)\right|\cdots\sup_{|\eta|\geq\frac{l}{L}\delta}\left|\hat{f}\Big(\frac{\sigma_{n}}{\bar{\sigma}}\eta\Big)\right|\\ \int_{|\eta|\geq\frac{l}{L}\delta}\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\right|\cdots\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta.$ (18)
@@
$\sup_{|\eta|\geq\frac{l}{L}\delta}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|=\sup_{|\eta^{\prime}|\frac{\bar{\sigma}}{\sigma_{j}}\geq\frac{l}{L}\delta}|\hat{f}(\eta^{\prime})|\leq\sup_{|\eta^{\prime}|\geq\frac{l^{2}}{L^{2}}\delta}|\hat{f}(\eta^{\prime})|\leq\kappa(\delta)$ (19)
@@
$\begin{split}\int_{|\eta|\geq\frac{l}{L}\delta}&\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\right|\cdots\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta\\ &\leq\bigg(\int_{\mathbb{R}^{d}}\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg)^{1/p^{\prime}}\cdots\bigg(\int_{\mathbb{R}^{d}}\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg)^{1/p^{\prime}}\\ &=\Big(\frac{\bar{\sigma}}{\sigma_{1}}\Big)^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\cdots\Big(\frac{\bar{\sigma}}{\sigma_{p^{\prime}}}\Big)^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\leq\Big(\frac{L}{l}\Big)^{d}\|\hat{f}\|_{p^{\prime}}^{p^{\prime}}\leq\Big(\frac{L}{l}\Big)^{d}C_{p}\|f\|_{p}^{p^{\prime}},\end{split}$ (20)
@@
We proceed analogously for $T_{2}$, obtaining a very similar bound. Since $\kappa(\delta)<1$, $\kappa(\delta)^{n-p^{\prime}}$ decays very fast in $n$, thus for $n$ large enough (depending on $p$), there exists $C_{1}>0$ such that
@@
$\left|\hat{G}\Big({\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}}\Big)\right|=e^{-\frac{|\xi|^{2}}{2n}\frac{\sigma_{k}^{2}}{\bar{\sigma}^{2}}}\leq e^{-\frac{|\xi|^{2}}{2n}(\frac{l}{L})^{2}}.$
@@
$\begin{split}&\frac{\left|\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}(\xi)\right|}{|\xi|^{2+s}}\\ &=\frac{1}{\left|\xi\right|^{2+s}}\Big[\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=1}^{j-1}\hat{f}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=j+1}^{n}\hat{G}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\Big]\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|e^{-\frac{|\xi|^{2}}{4n}(j-1)(\frac{l}{L})^{2}}e^{-\frac{|\xi|^{2}}{2n}(n-j)(\frac{l}{L})^{2}}\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{j=1}^{n}\left|\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}\frac{\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|}{\left|\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}}e^{-\frac{|\xi|^{2}}{4}\frac{n-1}{n}(\frac{l}{L})^{2}}\\ &\leq\frac{1}{n^{\frac{2+s}{2}}}\Big(\frac{L}{l}\Big)^{2+s}n\sup_{\eta^{\prime}}\frac{|\hat{f}(\eta^{\prime})-\hat{G}(\eta^{\prime})|}{\left|\eta^{\prime}\right|^{2+s}}e^{-\frac{\left|\xi\right|^{2}}{8}(\frac{l}{L})^{2}}.\end{split}$
@@
$\begin{split}\sup_{\eta^{\prime}}\frac{|\hat{f}(\eta^{\prime})-\hat{G}(\eta^{\prime})|}{\left|\eta^{\prime}\right|^{2+s}}&\leq\sup_{\eta^{\prime}}\left(\frac{|\hat{f}(\eta^{\prime})-\hat{f}(0)|}{\left|\eta^{\prime}\right|^{2+s}}+\frac{|\hat{G}(\eta^{\prime})-\hat{G}(0)|}{\left|\eta^{\prime}\right|^{2+s}}\right)\\ &\leq C_{1}(\rho_{2+s}+M_{2+s}(G)).\end{split}$
@@ 3.4 Positivity of the solution
Let $X_{m}$ be a Poisson random variable of parameter $m$, so that
$s_{m}=\mathbb{P}(m\leq X_{m}\leq 2m).$
For a fixed $m$, using that $\mathbb{P}(X_{m}\geq m)\geq 1/2$ and the Chernoff-type bound $\mathbb{P}(X_{m}\geq x)\leq e^{-m}(em)^{x}/x^{x}$ (for $x>m$), we have that
$s_{m}=\mathbb{P}(X_{m}\geq m)-\mathbb{P}(X_{m}\geq 2m+1)\geq\frac{1}{2}-e^{m+1}\Big(\frac{m}{2m+1}\Big)^{2m+1}:=b_{m}\geq b_{1}=:C_{L},$
since $b_{m}$ is increasing for $m\geq 1$. To prove that $s_{m}\to 1/2$, let $Y_{i},$ $i=1,\dots,m$ be Poisson-distributed random variables with parameter $1$ so that $X_{m}=\sum_{i=1}^{m}Y_{i}$. We have
$\begin{split}&e^{-m}\sum_{n=m}^{2m}\frac{m^{n}}{n!}=e^{-m}\sum_{n=0}^{2m}\frac{m^{n}}{n!}-e^{-m}\sum_{n=0}^{m-1}\frac{m^{n}}{n!}\\ &=\mathbb{P}\Big(\sum_{i=1}^{m}Y_{i}-m\leq{m}\Big)-\mathbb{P}\Big(\frac{1}{\sqrt{m}}\big({\sum_{i=1}^{m}Y_{i}-m}\big)\leq-\frac{1}{\sqrt{m}}\Big)\ \rightarrow 1-\frac{1}{2}=\frac{1}{2},\end{split}$
@@
Theorem 3.5 gives the confining Lyapunov conditions in both case $(i)$ and $(ii)$, concluding the proof ∎